OpenStudy (anonymous):

Limt x-x^2/x+4-sqrt(x^2+8) X->-1

OpenStudy (anonymous):

\[\frac{x-x^2}{-\sqrt{x^2+8}+x+4} \]Is your expression the one above?

OpenStudy (anonymous):

Yes correct

OpenStudy (anonymous):

\[\frac{ 1-x^2 }{ (x+4)-Sqrt(x^2+8) }\]

OpenStudy (anonymous):

So is it (x - x²) in the numerator, or (1 - x²)

OpenStudy (anonymous):